There are two environments for quotations: quote and quotation. quote is used either for a short quotation or for a sequence of short quotations separated by blank lines:

US presidents ... pithy remarks:
\begin{quote}
The buck stops here.
I am not a crook.
\end{quote}

US presidents have been known for their pithy remarks:

The buck stops here. I am not a crook.

Use the quotation environment for quotations that consist of more than one paragraph. Paragraphs in the input are separated by blank lines as usual:

Here is some advice to remember:
\begin{quotation}
Environments for making
...other things as well.
Many problems
...environments.
\end{quotation}

Here is some advice to remember:

Environments for making quotations can be used for other things as well.

Many problems can be solved by novel applications of existing environments.